A Bibliography of Supercomputing ’2000

Nelson H. F. Beebe University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 1.04

Abstract

This bibliography records articles presented at the Supercomputing ’2000 conference.

Title word cross-reference

3 [KLSK00, RT00].

/Mflops/s [ABE00].

1.34 [NSK^{+}00].

2 [GLT00, TRH00]. 2000 [JW00, Kra00, Lif00, MT00].

5 [TRH00].

6 [MFK00].

92/ c [ABE00].

Accelerate [FG00]. Access [BMLP^{+}00, Gre00, CDOS00, OD00]. Accurate [HD00]. across [CDS00]. ACTS [Dru00]. Adaptation [PK00]. Adaptive [CCD^{+}00, LTB00, OHS00, SKK00, SSOB00]. Adding [CB00]. Address [Wal00]. Adequate [OMS^{+}00]. Administering [Jon00b]. Advanced [GLT00]. Advancing [Kai00]. Advantages [Ede00, LBH00]. Aggregation [TT00]. Agreements [ZK00]. Algebra [DDS00]. Algorithm [SKK00]. Algorithms [DDS00, OHS00, PHBM00]. Alpha [Mor00a]. Analysis [AA00, BNPH00, HW00, LCM^{+}00, MGM00, MKT00, PKMB00, Ple00, Sol00, TWGS00]. Any [Gre00]. Anytime [FEL^{+}00]. Anywhere [FEL^{+}00, Gre00]. AppLeS [COBW00]. Appliances [Woo00].

Application [Bro00, BDG^{+}00, Chr00, Gre00, GKKS00, GMM00, GTH00, Lom00, LM00, ML00, May00, SBG00, TBW00]. Application-Driven [Chr00]. Application-level [SBG00, TBW00]. Applications [FG00, Gri00, GRP00, HW00, Mad00, MT00, SSOB00, SBG00, TWGS00, VdS00]. APPMAP [GTH00]. Approach [HRC^{+}00, HD00, KAF^{+}00]. Architectural [HLMR00]. Architecture [FG00]. Architectures [CDS00]. Area [AKF00, CDOS00, MC00, OD00, SBG00]. Arithmetics [HD00]. Art [Arm00]. ASCI [San00]. Assembly [Ist00]. Asynchronous [TBW00]. Audio [CST^{+}00]. Automatic [LF00]. Automatically [Irw00, VFD00]. Autonomous [HB00a]. Awards [Hae00, McG00].

Background [Hen00a]. Balancing [LTB00, SKK00]. Bandwidth [BNPH00, KS00, Par00]. Barriers [ZVD00]. Based [LGMZ00, MKT00, SSK^{+}00a, VT00, PFH^{+}00, SMAF00, TBW00]. Basic [Kea00]. Batch [Jon00a, Mat00a, PK00]. Be [WBKG00, CB00]. Beams [QRH00]. Benchmark [CDS00, WOK^{+}00]. Benchmarks [AA00, HCZ00, CE00]. Beneficial [CB00]. Benefits [Ede00]. Billionaire [WBKG00]. Biology [ZSS^{+}00]. Biomechanical [WFG^{+}00]. Biomolecular [BPK00, Ger00]. Black [MFK00]. Bleeding [Lin00]. block [LBH00]. Blue [Den00, Tur00]. Blurring [Fer00]. BOF [Bre00, Cat00, Dru00, Fis00, Irw00, Jon00a, Jon00b, Kai00, Lif00, Mas00, Mat00c, Mat00b, Mor00b, San00, Str00, dS00]. Bottlenecks [BH00]. Boundaries [Fer00]. Brain [WFG^{+}00]. Breaking [Mar00]. Broker [VT00]. Buffer [LF00]. Building [Kau00, LM00, ML00].

Caches [BTL^{+}00, BSL^{+}00]. Caching [Voe00]. Candidate [BMLP^{+}00]. Capabilities [Nos00]. Capital [WBKG00]. Cardiac [PHB^{+}00]. Case [Fah00, SBG00, TAK^{+}00]. Center [ACM00, MFK00]. CFD [GKKS00, HML^{+}00, RH00]. CG [TAK^{+}00]. Challenge [Fre00]. Challenges [Rat00]. Characterization [AA00]. Charged [QRH00]. Charged-Particle [QRH00]. Circulation [LBH00]. ClassAd [VT00]. Climate [CKF^{+}00]. Cluster [ABE00, AKF00, BHP00, BBD00, Bec00, HML^{+}00, Mat00b, MOK^{+}00, Ste00b, YL00]. Clustered [PNMB00]. Clusters [Bjo00, Bre00, FG00, HLMR00, Mor00a, PKMB00, Rho00]. Cockpit [Bec00]. Code [AMPS00, CB00]. Codes [AJ00, LRSW00]. Coherency [AJ00]. Collaboration [CDOS00, OD00]. Collaborator [Kee00]. Collection [BNPH00]. Collective [VFD00]. Collider [MM00]. Collisions [QRH00]. Commercial [Lom00]. Commodity [Bjo00, PFH^{+}00]. Commodity-based [PFH^{+}00]. Communication [TSH^{+}00, TRH00]. Communications [VFD00]. Community [Ter00]. Comparative [CDS00]. Comparison [SSOB00]. Compiler [GMM00, Mil00]. Compilers [Wol00]. Components [LRSW00]. Composites [KLSK00]. Computational [FT00, Kea00, SFP^{+}00, SLJ^{+}00, ZSS^{+}00]. Computations [RT00]. Compute [Ede00]. Computer [NSK^{+}00, PHB^{+}00, San00]. Computers [BHP00, Hen00a]. Computing [ACM00, AKF00, Arm00, BBD00, Bjo00, BESW00, CS00, Chr00, Ede00, FEL^{+}00, Ger00, GSC^{+}00, GMM00, HSC00, Ist00, JGT^{+}00, Kai00, KAF^{+}00, Owe00, Lif00, Mad00, Mat00b, Mil00, MKT00, Mor00b, MOK^{+}00, OMS^{+}00, PNK00, PNMB00, Rat00, SSK^{+}00a, Ste00b, SJ00a, SJ00b, SMAF00, Wri00, ZSS^{+}00, dS00, Mat00b]. Concept [Ple00]. Concept-to-Production [Ple00]. Concurrent [BG00]. Consistent [QRH00]. Consortium [Kai00]. Constraints [RH00]. contemporary [Mar00]. Convention [ACM00]. Cooperative [Voe00]. CORBA [LRSW00, TBW00]. CORBA-based [TBW00]. Core [HB00a]. Corner [PSC^{+}00]. Cost [Bjo00, HML^{+}00, Mor00b]. COTS [Ste00b]. Coulomb [QRH00]. Counters [BDG^{+}00]. CPlant [Lin00]. Cross [BDG^{+}00]. Cross-Platform [BDG^{+}00]. Current [BBD00]. Curvilinear [LBH00]. Cyber [Ist00]. Cyber-Pharmaceutical [Ist00].

D [KLSK00, RT00]. Dallas [ACM00]. Data [AJ00, BMLP^{+}00, BTL^{+}00, BSL^{+}00, CF00, GK00, HC00, HCZ00, JGT^{+}00, LGMZ00, MC00, NPP^{+}00, NT00, SBG00]. Database [BMLP^{+}00, Guz00, NT00, TBW00]. Databases [BNPH00]. Dataset [SSK^{+}00a]. Datasets [SSK00b]. DataSpace [GRC^{+}00]. Date [CKF^{+}00]. DB [Guz00]. Deep [Tur00]. Definition [GLP00, PGM^{+}00]. Deformation [WFG^{+}00]. Delegation [SNT00]. Delivering [Pas00]. Demonstrations [MT00]. Demos [LM00, ML00]. Design [KLSK00, PKMB00, Ste00a]. Designs [OHS00]. Developers [May00]. Development [Dru00]. Developments [Mat00c]. Directions [DDS00]. Directives [BKO00]. Discovery [Kea00]. Discrete [Hen00b]. Displays [HBEH00, PS00]. Distributed [BTL^{+}00, BSL^{+}00, Bro00, CS00, Ede00, HBEH00, Mad00, SMAF00, SNI00, TBW00, WBW^{+}00, ZK00]. Distribution [NPP^{+}00]. DMP [Lom00]. DoD [Cam00]. Dose [Spa00]. Driven [Chr00]. Drug [BMLP^{+}00, Kea00]. DSM [AJ00]. Dual [Fah00, LBH00]. Dual-Level [LBH00]. DYNA [Chr00]. Dynamic [ABG^{+}00, BMLP^{+}00, LTB00, LCM^{+}00, VdS00]. Dynamic-Mesh [ABG^{+}00].

Dynamics [BPK00, NSK^{+}00].

EARTH [TAK^{+}00, Tan00]. Edge [Lin00]. Education [HB00b, Jam00]. Effective [Bjo00, SJ00a, SJ00b]. Efficient [BESW00, CS00]. Electrophysiology [PHB^{+}00]. Element [Hen00b]. Emerging [BBD00]. Enable [BTL^{+}00, BSL^{+}00, Gre00]. Enabled [CST^{+}00]. Enabling [Cam00, KAF^{+}00]. End [Ple00]. Enforcing [ZK00]. Engine [Guz00]. Engineering [JGT^{+}00, RH00]. Engines [KKC00]. Enotebook [GBG^{+}00]. Environment [CS00, PKF^{+}00, PNMB00, SF00, VT00]. Environments [TSH^{+}00]. Era [BGG^{+}00]. eSCape [JW00, MT00]. eSCaped [GZ00]. eSCapes [ZVD00]. ESP [WOK^{+}00]. Evaluation [GSC^{+}00, HLMR00]. Event [BNPH00]. Events [MM00]. Evolutionary [TAK^{+}00]. executable [PKF^{+}00]. Execution [SF00]. Existing [CB00]. Experiences [Mas00]. Exploration [SSK^{+}00a]. Expressing [ZK00]. Extending [BCC^{+}00, Mat00a, Tho00].

Failure [Bro00, FT00]. Farms [Ede00]. Fault [SNI00, YL00]. Fault-Tolerant [SNI00]. Features [GLT00, Mil00]. File [NT00, PNMB00]. Fill [Rho00]. Fine [HCZ00, TAK^{+}00]. Fine-Grained [HCZ00, TAK^{+}00]. First [Ist00]. Flexibility [Pas00]. Flow [CCD^{+}00]. Flows [ABG^{+}00]. Fluid [CCD^{+}00]. Force [Mat00b]. Formats [BESW00]. Forum [Cat00]. fragmentation [Mar00]. Framework [AMPS00, HC00, LCM^{+}00, WBW^{+}00]. FSL [Lin00]. Fujitsu [Miu00]. Future [Fis00, Mat00c].

Galactic [MFK00]. Games [SK00]. Gara [Tho00]. Gene [Den00]. General [PNK00]. Generation [Fre00, Owe00, Mad00, MKT00, WBW^{+}00].

Generator [LRSW00]. Genome [Mes00]. Gigabit [GLP00, PGM^{+}00]. Gigabit/sec [GLP00, PGM^{+}00]. Gigabyte [PNMB00]. GigaNet [HLMR00]. Global [Cat00]. Globus [BNPH00, SNT00, VT00]. GPFS [Jon00b]. GQ [RFG^{+}00]. Grained [HCZ00, TAK^{+}00]. GRAPE [MFK00]. GRAPE-6 [MFK00]. Grid [COBW00, Cat00, Owe00, LBH00, SSK^{+}00a, TWGS00, BNPH00, CDOS00, FEL^{+}00, OD00, Tho00]. GridPort [MT00]. Grids [FT00, JGT^{+}00, PHBM00, SFP^{+}00, SLJ^{+}00]. GridSearcher [MS00]. Group [dS00]. Guided [WFG^{+}00].

Hardware [AJ00, BDG^{+}00, BH00]. HDF5 [CF00]. Hearing [ZVD00]. Heavy [MM00]. Heterogeneous [CS00, LGMZ00, SMAF00, TSH^{+}00]. High [ACM00, AMPS00, BTL^{+}00, BSL^{+}00, Bjo00, BESW00, CCD^{+}00, CF00, CDOS00, Chr00, DDS00, FT00, Fis00, FEL^{+}00, Fre00, GK00, GMM00, Guz00, HML^{+}00, Irw00, KAF^{+}00, KS00, Owe00, LRSW00, Lif00, MC00, Mil00, MKT00, Miu00, Mor00a, Mor00b, NT00, Nos00, OD00, OMS^{+}00, PS00, Par00, PNK00, PKMB00, Ple00, SBG00, Ste00b, TSH^{+}00, ZSS^{+}00, GLP00, PGM^{+}00]. High-Cost [HML^{+}00]. High-level [AMPS00]. High-Performance [Bjo00, FT00, KAF^{+}00]. High-resolution [CDOS00, OD00, PS00]. High-Speed [BTL^{+}00, DDS00, BSL^{+}00]. Highly [Miu00]. HiMAP [GRP00]. History [SSK^{+}00a]. Holes [MFK00]. Hot [FG00]. HPC [Cam00, KLSK00, Mar00, Pas00, SK00]. Human [Mes00]. Hybrid [Hen00b].

I/O [May00, NT00]. IBM [CE00, dS00]. Idle [HSC00]. IEEE [Mat00b]. II [EM00b, ML00, MKT00, SJ00b]. Image [WFG^{+}00]. Imperfectly [AMP00]. Imperfectly-nested [AMP00]. Implementation [TBW00, TRH00, VT00].

Implementations [BM00]. Implementing [Woo00]. Improving [HCZ00, LTB00]. Incorporating [AKF00]. Industry [Rat00]. Inference [SWCM00]. InfiniBand [Ano00]. Infosec [Spa00]. Infrastructure [BDG^{+}00, CKF^{+}00]. Infrastructures [Woo00]. Innovative [HRC^{+}00]. Insight [Wri00]. Integrated [PKF^{+}00]. Integrating [FEL^{+}00, NT00]. Integration [LGMZ00]. Intel [Wol00]. Intelligent [Woo00]. Intensive [SBG00]. Interactive [SSK^{+}00a]. Interconnects [Fis00, HLMR00]. Interface [GLT00]. Interfaces [ABG^{+}00]. Internet [Arm00, BGG^{+}00, HB00a, Mad00, Ste00a]. Interoperability [AKF00]. Intrepid [Kee00]. Introduction [GK00, SJ00a, SJ00b]. Ion [MM00]. IP [BGG^{+}00, GLP00, PGM^{+}00]. IPV6 [HDK^{+}00]. Irregular [HCZ00]. Isolate [BH00]. Issues [AKF00]. IU [GBG^{+}00].

Japan [Tan00]. Java

[BKO00, GMM00, LRSW00]. Java/CORBA [LRSW00]. Job [Mat00a, SF00]. Juniper [Jam00].

KAI [Wol00]. Keynote [Wal00].

Lagrangian [ABG^{+}00]. LAN [Fer00]. Landing [TAK^{+}00]. Langevin [QRH00]. Language [GMM00]. Languages [CDS00]. Large [ABE00, BMLP^{+}00, BPK00, Ede00, Ger00, GRP00, HC00, HW00, Lom00, MOK^{+}00, SSK00b]. Large-Scale [Ede00, HW00, GRP00]. Leads [Mad00]. Legacy [LRSW00]. Legion [Gri00]. Level [COBW00, Fah00, LBH00, AMPS00, SBG00, TBW00]. Levels [LCS00]. Liberating [Ste00a]. Library [PNK00]. like [BKO00]. Linear [DDS00]. Linux [Lin00]. Liquid [KKC00]. Literacy [Mor00b]. Load [LTB00, SKK00]. Load-balancing [SKK00]. Looking [SFP^{+}00]. Loop [AMP00]. Lots [Wri00]. Low [HML^{+}00, Mor00b]. Low-Cost [HML^{+}00]. LS [Chr00]. LS-DYNA [Chr00].

M1 [HW00]. M10 [TWGS00]. M2 [May00]. M3 [HC00]. M4 [ZSS^{+}00]. M5A [LM00]. M5B [ML00]. M6A [EM00a]. M6B [EM00b]. M7 [DDS00]. M8 [BHP00]. M9 [BBD00]. Machines [BCC^{+}00, TT00]. MAN [Fer00]. MAN/LAN/WAN [Fer00]. Management [AKF00, Ano00, CKF^{+}00, Gam00, NT00]. Managing [SSK00b]. Manipulating [SSK00b]. Massively [MKT00, TT00]. Matching [SSK^{+}00a]. Matching-Grid [SSK^{+}00a]. Matrix [AMPS00]. MD [HRC^{+}00]. MDM [NSK^{+}00]. Meaningful [PSC^{+}00]. Mechanisms [CS00]. Megacomputers [SCF^{+}00]. Memory [BH00, Hen00b, HCZ00, SNI00]. Mesh [ABG^{+}00, CCD^{+}00, GKKS00, Owe00, LTB00, MKT00, MKT00]. Message [GLT00, Hen00b, RFG^{+}00]. Message-Passing [GLT00, Hen00b]. Metacomputing [MSB^{+}00]. Method [ABG^{+}00]. Methods [HC00, TBW00]. Mflops [ABE00]. MG [CDS00]. MicroGrid [SLJ^{+}00]. Middleware [BNPH00, COBW00, GZ00, TSH^{+}00]. Migration [CS00]. Millennium [CF00]. Mining [GK00]. Minutes [GTH00]. Model [LGMZ00]. Model-Based [LGMZ00]. Modeling [CKF^{+}00, GKKS00, Hen00b, LBH00, SLJ^{+}00]. Models [LCS00, SSOB00]. Molecular [BPK00, HRC^{+}00, NSK^{+}00]. Monitoring [Ple00, SMAF00]. Monitors [BH00]. Moving [Lom00]. MPI [BM00, CE00, CB00, GLT00, TRH00, VdS00]. MPI-2 [GLT00, TRH00]. MPICH [RFG^{+}00]. MPICH-GQ [RFG^{+}00]. Multi [LBH00]. Multi-block [LBH00]. Multidisciplinary [GRP00]. Multigrid [TT00]. Multithreading [TAK^{+}00]. Myrinet [HLMR00].

NaCl [NSK^{+}00]. NAS [CE00, CDS00]. NEC [TRH00]. Necessary [NPP^{+}00]. Needs [OMS^{+}00]. nested [AMP00]. Nests [AMP00]. Network [ABE00, AKF00, BTL^{+}00, BSL^{+}00, Gam00, Kee00, MC00, Ple00, SBG00, SMAF00, TSH^{+}00]. Network-Computing [AKF00]. Networking [ACM00]. Networks [Fre00, Irw00, Jam00, SBG00]. Neural [ABE00]. Neural-Network [ABE00]. Neuroscience [LGMZ00]. Neurosurgery [WFG^{+}00]. Next [FG00, Fre00]. November [ACM00]. NUMA [BCC^{+}00]. Numbers [Wri00]. Numerical [DDS00, GMM00].

O [May00, NT00]. Object

[BNPH00, LCM^{+}00, SF00, TBW00]. Object-Oriented [LCM^{+}00, SF00]. Objectivity [Guz00]. Objectivity/DB [Guz00]. Ocean [LBH00]. One [TRH00]. One-Sided [TRH00]. op [Ger00]. Open [BGG^{+}00, Ter00]. OpenMP [BKO00, CB00, NPP^{+}00, BCC^{+}00, CE00, EM00a, EM00b, Mat00c]. OpenMP-like [BKO00]. Optimization [Wol00]. Optimizations [RT00]. Oriented [LCM^{+}00, SF00]. Origin2000 [SSOB00]. Overview [Owe00].

Panel [BGG^{+}00, OMS^{+}00, PSC^{+}00, SFP^{+}00, SCF^{+}00, WBKG00]. Papers [McG00]. Papers/Awards [McG00]. Paradigm [Ist00]. Parallel [ABG^{+}00, BHP00, Bro00, BKO00, CDS00, EM00a, EM00b, GRP00, HSC00, Kai00, KKC00, May00, MGM00, MKT00, NT00, PKF^{+}00, PHBM00, RH00, SWCM00, SJ00a, SJ00b, TT00, WBW^{+}00]. Parallel-Based [MKT00]. Parallel/Distributed [Bro00]. Parallelism [Fah00, Hen00b, LBH00, LCS00]. Parallelization [Wol00]. Parameter [COBW00]. Part [EM00a, EM00b, Owe00, LM00, ML00, MKT00, SJ00a, SJ00b].

Particle [BNPH00, LCS00, QRH00]. Passing [GLT00, Hen00b, RFG^{+}00]. Path [TAK^{+}00]. PBS [Jon00a, Tho00]. Peek [Fis00]. Peformance [Lif00]. Performance [ACM00, AKF00, BMLP^{+}00, Bjo00, BDG^{+}00, BH00, BESW00, CCD^{+}00, CF00, Chr00, FG00, FT00, FEL^{+}00, Fre00, GKKS00, GK00, GMM00, GTH00, Guz00, Hen00b, HW00, HLMR00, Irw00, KAF^{+}00, Owe00, LRSW00, MC00, Mil00, MGM00, MKT00, Miu00, Mor00a, Mor00b, NT00, Nos00, OMS^{+}00, Par00, PNK00, PKMB00, Ple00, Sol00, Ste00b, TSH^{+}00, TWGS00, Voe00, Wes00, WBW^{+}00, ZSS^{+}00]. Perspective [Gri00]. Peta [Ger00]. Peta-op [Ger00]. Petaflops [PSC^{+}00, Wal00]. Pharmaceutical [BMLP^{+}00, Ist00]. Phylogenetic [SWCM00]. Physics [BNPH00]. PIII [ABE00]. Plans [Hen00a, Nos00]. Platform [BDG^{+}00, Gre00]. Platforms [HD00]. Playback [PS00]. PM2 [TSH^{+}00]. Policy [SNT00]. Portable [PNK00, Jon00a, Mat00a]. Portal [FEL^{+}00]. Portals [KAF^{+}00]. Power [Wes00]. Predict [GTH00]. Prediction [AJ00, HW00]. Preemptive [Mat00a]. Problem [SFP^{+}00]. Problems [GRP00]. Process [CS00, GRP00]. Processing [RH00]. Processors [Bjo00, CCD^{+}00]. Production [Ple00]. Program [HB00b, Mas00]. Programming [BHP00, BG00, Bro00, BKO00, EM00a, EM00b, PKF^{+}00, SSOB00]. Programs [MGM00, RFG^{+}00]. Project [GRC^{+}00, Tan00]. Promoting [Mor00b]. Prospectus [San00]. Protocol [OMS^{+}00]. Protocols [GSC^{+}00]. Proxies [SNT00]. Proxy [Voe00]. PSockets [SBG00]. Pthreads [BG00]. pump [KKC00]. PUNCH [FEL^{+}00, KAF^{+}00]. Purpose [NSK^{+}00].

QoS [CST^{+}00]. Quality [RFG^{+}00, Tho00].

Quality-of-Service [RFG^{+}00].

Radiation [PHBM00]. RAIN [Bru00]. Rainfinity [Bru00]. Randomization [PK00]. Rapid [KLSK00]. Reactive [CCD^{+}00]. Real [Sol00, WFG^{+}00]. Real-Time [WFG^{+}00]. Real-World [Sol00]. Realization [KLSK00]. Realizing [YL00]. Refinement [CCD^{+}00, LTB00]. Relativistic [MM00]. Remote [BTL^{+}00, BSL^{+}00, Gam00]. Rendering [HBEH00]. Reordering [HCZ00]. Requirements [GSC^{+}00]. Research [Bro00, Cam00, CKF^{+}00, DDS00, Jam00, Kea00, Mas00]. Reservoir [SSK^{+}00a]. Resilience [YL00]. resolution [CDOS00, OD00, PS00]. Resource [ZK00]. Resources [MGM00]. Restricted [SNT00]. RMI [GSC^{+}00]. Rocket [KKC00]. Routers [Jam00]. Routing [Fer00].

s [ABE00]. S1 [GLT00]. S10 [PFH^{+}00]. S2 [GK00]. S3 [PKMB00]. S4 [GMM00]. S5 [MGM00]. S6A [Owe00]. S6B [MKT00]. S7A [SJ00a]. S7B [SJ00b]. S8 [SSK00b]. S9 [BG00]. SAN [Woo00]. Satisfying [RH00]. SC2000 [ACM00, HB00b, McG00, SK00]. Scalable [BDG^{+}00, BPK00, CDOS00, HBEH00, Miu00, OHS00, OD00, PFH^{+}00, SMAF00, SNI00]. Scale [ABE00, Ede00, Ger00, HW00, HLMR00, MOK^{+}00, Voe00, GRP00]. Scheduled [HSC00]. Schedulers [PK00]. Scheduling [Mat00a]. SCICOMP [dS00]. Science [CF00, JGT^{+}00, San00]. Scientific [AJ00, GSC^{+}00, HW00, NT00, RT00, SSK00b, SKK00, SLJ^{+}00, Wri00, dS00]. SCinet [Kra00]. SDSC [MT00]. sec [GLP00, PGM^{+}00]. Second [PNMB00]. Self [QRH00]. Self-Consistent [QRH00]. Server [YL00]. Servers [HLMR00, Miu00]. Service [RFG^{+}00, Tho00]. Services [FEL^{+}00]. Session [Hae00]. Shared [Hen00b, HCZ00, SNI00]. Shared-Memory [Hen00b, HCZ00]. Sharing [ZK00]. SHMEM [PNK00]. Sided [TRH00, BM00]. Sign [ZVD00]. Similarity [AA00]. Simplicity [Wes00]. Simulation [ABG^{+}00, Ger00, HRC^{+}00, MFK00, Mar00, NSK^{+}00, QRH00, SSK^{+}00a, San00, WFG^{+}00]. Simulations [CCD^{+}00, KKC00, MOK^{+}00, PHB^{+}00, SKK00]. Simulator [Tan00]. Single [BM00]. Single-sided [BM00]. Small [HLMR00, Spa00]. Small-Scale [HLMR00]. Smoothed [TT00]. SMP [HLMR00, Lom00]. SNMP [SMAF00]. SNMP-based [SMAF00]. Socket [LF00]. Software [FG00, LCM^{+}00, VdS00]. Solution [CF00, SFP^{+}00]. Solutions [Ano00, GMM00, Ple00]. Source [BGG^{+}00, Ter00]. Sources [LGMZ00]. SP [dS00, CE00]. Sparse [AMPS00]. SPECcpu95 [AA00]. Special [NSK^{+}00]. Special-Purpose [NSK^{+}00]. Specification [SNT00]. Specifications [AMPS00]. Speculation [PK00]. Speech [ZVD00]. Speech-to-Sign [ZVD00]. Speech-to-Text [ZVD00]. Speed [BTL^{+}00, DDS00, Fis00, SBG00, BSL^{+}00]. SRC [Hen00a]. Staggering [Wes00]. stagnation [Mar00]. Standards [LM00, ML00]. Static [LCM^{+}00]. Statistical [OHS00]. Status [Fis00, Hen00a, MC00, Tan00]. Stochastic [Mar00]. Storage [Gam00, Nos00, VT00]. Strategies [Chr00, TT00]. Striping [SBG00]. Structured [Owe00]. Study [CDS00, Fah00, TAK^{+}00]. STWAVE [Fah00]. Success [Bro00]. SUN [BM00]. Supercomputers [KS00, Miu00, Str00]. Supercomputing [Kau00, Par00]. Support [NT00, SSK00b]. Sweep [COBW00]. Switching [Fer00]. SX [TRH00]. SX-5 [TRH00]. Synthesis [AMPS00]. System [HSC00, HB00a, Jon00a, Kau00, MOK^{+}00, Nos00, SSK00b, SMAF00, WOK^{+}00, Mat00a]. Systems [AKF00, BPK00, Ste00b, WBW^{+}00].

Tarantella [Gre00]. Task [Mat00b]. TCP [FT00, Irw00, LF00, OMS^{+}00]. Teaming [Wes00]. Teams [Ste00a]. Technical [McG00]. Techniques [Owe00, LTB00]. Technologies [FEL^{+}00, HC00, KAF^{+}00]. Technology [San00, Wes00]. Telemiscroscopy [HDK^{+}00]. Teleportation [CST^{+}00]. Template [COBW00]. Templates [LCM^{+}00]. TeraOPS [Kau00]. Testing [VdS00]. Text [ZVD00]. Tflops [MFK00, NSK^{+}00]. Third [Mad00]. Thirsty [BNPH00]. Thousands [CCD^{+}00]. Three [SSOB00]. Tiled [PS00]. Tiling [AMP00, RT00]. Time [MC00, WFG^{+}00]. Time-Varying [MC00]. Today [KS00]. Tolerant [SNI00]. Tomorrow [KS00]. too [Wri00]. Tool [LCM^{+}00, PKF^{+}00, SLJ^{+}00]. Toolkit [Dru00, MT00]. Tools [Kai00, LM00, ML00, MGM00, SSK00b, Wol00]. TOP500 [Str00]. Topology [HB00a]. Trace [WBW^{+}00]. Tracer [LCS00]. Training [ABE00]. Transfer [PNMB00]. Transport [LCS00, PHBM00]. Transportation [Rat00]. Travel [Rat00]. Trends [BBD00]. Tuned [VFD00]. Tuning [BDG^{+}00, GKKS00, Irw00, LF00, MGM00, TWGS00]. Turbo [KKC00]. Turbo-pump [KKC00]. TurboLinux [Bec00]. Tutorial [BHP00, BBD00, BG00, DDS00, EM00a, EM00b, GLT00, GK00, GMM00, HC00, HW00, Owe00, LM00, ML00, May00, MGM00, MKT00, PFH^{+}00, PKMB00, SSK00b, SJ00a, SJ00b, TWGS00, ZSS^{+}00]. TV [GLP00, PGM^{+}00]. Two [LCS00]. TX [ACM00].

Ultra [ABE00]. Ultra-Large-Scale

[ABE00]. Umpire [VdS00].

Undergraduates [Mas00]. Unified

[SKK00]. Unsteady [KKC00].

Unstructured [GKKS00, Owe00, PHBM00]. Unveiling [Mes00]. Update [Tur00]. USA [ACM00]. Use [Jam00]. User [COBW00, dS00]. User-Level [COBW00]. Users [Lif00]. Using [BTL^{+}00, BSL^{+}00, Bre00, BDG^{+}00, BH00, BNPH00, CCD^{+}00, GLT00, Jon00b, Woo00, SBG00]. Utilization [WOK^{+}00]. Utilizing [HSC00].

Varying [MC00]. Vendor [Rho00]. Venture [DW00, WBKG00]. versus [CE00]. VG [MOK^{+}00]. VI [FG00]. Via [Ste00a]. Village [DW00]. Visapult [BSL^{+}00]. Visual [HRC^{+}00, MOK^{+}00]. Visualization [BTL^{+}00, BSL^{+}00, Bre00, BNPH00, HC00, MC00, MM00, PS00, PFH^{+}00, WBW^{+}00]. Void [Rho00]. Volume [MC00]. Volumetric [MOK^{+}00, WFG^{+}00].

WAN [Fer00]. WANs [BSL^{+}00, BTL^{+}00]. Wants [WBKG00]. Way [GTH00]. Web [PKF^{+}00, FEL^{+}00, GZ00, Gre00, Voe00, YL00]. Web-executable [PKF^{+}00]. Web-Server [YL00]. Web100 [Irw00]. Who [WBKG00]. Wide [AKF00, CDOS00, MC00, MSB^{+}00, OD00, SBG00]. Wide-Area [AKF00, MC00]. Will [CB00]. Windows [Lif00]. Wire [BESW00]. Within [AKF00]. Workload [AA00]. Workstations [HSC00]. World [MSB^{+}00, Pas00, Sol00]. Woven [KLSK00]. Wrapper [LRSW00]. Wrapping [LRSW00].

XML [LM00, ML00].

Year [Wal00].

References

Almojel:2000:WCS

[AA00] Abdullah I. Almojel and Ali S.

AlSwayan. Workload characterization and similarity analysis of SPECcpu95 benchmarks. In ACM [ACM00], page 153.

                                                   URL http://www.sc2000.org/
                                                   proceedings/info/fp.pdf.

Aberdeen:2000:MUL

[ABE00] Douglas Aberdeen, Jonathan Baxter, and Robert Edwards. 92/ c/mflops/s, ultra-large-scale neural-network training on a PIII cluster. In ACM [ACM00], pages 66–67. URL http://

                                                   www.sc2000.org/proceedings/
                                                   techpapr/papers/pap255.pdf.

Antaki:2000:PDM

[ABG^{+}00] James F. Antaki, Guy E. Blelloch, Omar Ghattas, Ivan Malcevic, Gary L. Miller, and Noel J. Walkington. A parallel dynamicmesh Lagrangian method for simulation of flows with dynamic interfaces. In ACM [ACM00], page 58. URL http://

                                                   www.sc2000.org/proceedings/
                                                   techpapr/papers/pap289.pdf.

ACM:2000:SHP

[ACM00] ACM, editor. SC2000: High Performance Networking and Computing. Dallas Convention Center, Dallas, TX, USA, November 4–10, 2000. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. URL http://

                                                   www.sc2000.org/proceedings/
                                                   info/fp.pdf.

Acquaviva:2000:HPD

[AJ00] J. T. Acquaviva and W. Jalby. Hardware prediction for data coherency of scientific codes on DSM. In ACM [ACM00], page 65. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap209.pdf.

Adabala:2000:PII

[AKF00] Sumalatha Adabala, Nirav H.

Kapadia, and José A. B. Fortes. Performance and interoperability issues in incorporating cluster management systems within a wide-area network-computing. In ACM [ACM00], page 53. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap262.pdf.

Ahmed:2000:TIN

[AMP00] Nawaaz Ahmed, Nikolay Mateev, and Keshav Pingali. Tiling imperfectly-nested loop nests. In ACM [ACM00], page 60. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap140.pdf.

Ahmed:2000:FSM

[AMPS00] Nawaaz Ahmed, Nikolay Mateev, Keshav Pingali, and Paul Stodghill. A framework for sparse matrix code synthesis from highlevel specifications. In ACM [ACM00], page 74. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap195.pdf.

Anonymous:2000:IMS

[Ano00] Anonymous. InfiniBand management solutions. In ACM [ACM00], page 101. URL http:/

/www.sc2000.org/proceedings/
info/fp.pdf.

Armentrout:2000:AIC

[Arm00] Steve Armentrout. The art of Internet computing. In ACM [ACM00], page 101.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Baker:2000:TMC

[BBD00] Mark Baker, Rajkumar Buyya, and Jack Dongarra. Tutorial M9: Current and emerging trends in cluster computing. In ACM [ACM00], pages 23–24.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Bircsak:2000:EON

[BCC^{+}00] John Bircsak, Peter Craig, Rae-Lyn Crowell, Zarka Cvetanovic, Jonathan Harris, C. Alexander Nelson, and Carl D. Offner. Extending OpenMP for NUMA machines. In ACM [ACM00], pages 68–69. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap226.pdf.

Browne:2000:SCP

[BDG^{+}00] S. Browne, J. Dongarra, N. Garner, K. London, and P. Mucci. A scalable cross-platform infrastructure for application performance tuning using hardware counters. In ACM [ACM00], page 65. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap256.pdf.

Beckman:2000:TCC

[Bec00]  Peter H. Beckman.  Tur-

boLinux Cluster Cockpit. In ACM [ACM00], page 102.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Bustamante:2000:EWF

[BESW00] Fabian E. Bustamante, Greg Eisenhauer, Karsten Schwan, and Patrick Widener. Efficient wire formats for high performance computing. In ACM [ACM00], page 64. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap171.pdf.

Breshears:2000:TSC

[BG00] Clay P. Breshears and Henry A.

Gabb. Tutorial S9: Concurrent programming with Pthreads. In ACM [ACM00], page 17.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Borchers:2000:POS

[BGG^{+}00] Robert Borchers, Susan Graham, Richard Gabriel, Todd Needham, and José Muñoz. Panel: Open source: IP in the Internet era. In ACM [ACM00], pages 79–80.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Buck:2000:UHP

[BH00] Bryan R. Buck and Jeffrey K.

Hollingsworth. Using hardware performance monitors to isolate memory bottlenecks. In ACM [ACM00], pages 64–65.

URL http://www.sc2000.org/
proceedings/techpapr/papers/
pap197.pdf.

Bader:2000:TMP

[BHP00] David A. Bader, Bruce Hendrickson, and Steve Plimpton. Tutorial M8: Parallel programming for cluster computers. In ACM [ACM00], page 22. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Bjornson:2000:CEH

[Bjo00] Robert D. Bjornson. Cost effective high-performance computing on clusters of commodity processors. In ACM [ACM00], page 98.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Bull:2000:PPJ

[BKO00] J. Mark Bull, Mark E. Kambites, and Jan Obdrzalek. Parallel programming in Java with openMPlike directives. In ACM [ACM00], page 150. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Booth:2000:SSM

[BM00] S. Booth and E. Mourao. Single-sided MPI implementations for SUN MPI. In ACM [ACM00], page 46. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap182.pdf.

Ben-Miled:2000:DAP

[BMLP^{+}00] Zina Ben-Miled, Yang Liu, Dave Powers, Omran Bukhres, Michael Bem, Robert Jones, Robert Oppelt, and Samuel Milosevich. Data access performance in a large and dynamic pharmaceutical drug candidate database. In ACM [ACM00], pages 55–56. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap213.pdf.

Bunn:2000:BTP

[BNPH00] J. Bunn, H. Newman, J. Patton, and K. Holtman. Bandwidth thirsty particle physics event collection analysis and visualization using object databases and the Globus Grid middleware. In ACM [ACM00], page 118.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Brunner:2000:SMD

[BPK00] Robert K. Brunner, James C.

Phillips, and Laxmikant V. Kale. Scalable molecular dynamics for large biomolecular systems. In ACM [ACM00], page 67. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap271.pdf.

Breckenridge:2000:BUC

[Bre00] Arthurine Breckenridge. BOF:

Using clusters for visualization. In ACM [ACM00], page 105.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Browne:2000:PDP

[Bro00]  J. C. Browne.  Parallel/

distributed programming: Research success — application failure? In ACM [ACM00], pages 29–30. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Bruck:2000:RR

[Bru00] Jehoshua Bruck. From RAIN to Rainfinity. In ACM [ACM00], pages 37–38. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Bethel:2000:VUH

[BSL^{+}00] W. Bethel, J. Shalf, S. Lau, D. Gunter, J. Lee, B. Tierney, V. Beckner, J. Brandt, D. Evensky, H. Chen, G. Pavel, J. Olsen, and B. H. Bodtker. Visapult — using high-speed WANs and network data caches to enable remote and distributed visualization. In ACM [ACM00], pages 118–119. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Bethel:2000:UHS

[BTL^{+}00] Wes Bethel, Brian Tierney, Jason Lee, Dan Gunter, and Stephen Lau. Using highspeed WANs and network data caches to enable remote and distributed visualization. In ACM [ACM00], page 59. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap125.pdf.

Campbell:2000:EHR

[Cam00]  Tim Campbell.  Enabling

HPC research in DoD. In ACM [ACM00], page 102.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Catlett:2000:BGG

[Cat00] Charlie Catlett. BOF: Global Grid Forum. In ACM [ACM00], page 110. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Cartwright:2000:AOE

[CB00] Keith L. Cartwright and Joseph D. Blahovec. Adding openMP to an existing MPI code: Will it be beneficial? In ACM [ACM00], page 145. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Calder:2000:HPR

[CCD^{+}00] A. C. Calder, B. C. Curtis, L. J. Dursi, B. Fryxell, G. Henry, P. MacNeice, K. Olson, P. Ricker, R. Rosner, F. X. Timmes, H. M. Tufo, J. W. Turan, and M. Zingale. High performance reactive fluid flow simulations using adaptive mesh refinement on thousands of processors. In ACM [ACM00], pages 72–73. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap303.pdf.

Childers:2000:SHR

[CDOS00] Lisa Childers, Terry Disz, Bob Olson, and Rick Stevens. Scalable high-resolution wide area collaboration over the Access Grid. In ACM [ACM00], page 121. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Chamberlain:2000:CSN

[CDS00] Bradford L. Chamberlain, Steven J.

Deitz, and Lawrence Snyder. A comparative study of the NAS MG benchmark across parallel languages and architectures. In ACM [ACM00], page 67. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap199.pdf.

Cappello:2000:MVM

[CE00] Franck Cappello and Daniel Etiemble. MPI versus MPI+OpenMP on the IBM SP for the NAS Benchmarks. In ACM [ACM00], page 51. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap214.pdf.

Cheng:2000:HHP

[CF00] Albert Cheng and Michael Folk. HDF5: High performance science data solution for the new millennium. In ACM [ACM00], page 149. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Christon:2000:LDA

[Chr00] Mark Christon. LS-DYNA — application-driven strategies for high performance computing. In ACM [ACM00], pages 42–43.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Chervenak:2000:DMI

[CKF^{+}00] A. Chervenak, C. Kesselman, I. Foster, S. Tuecke, W. Allcock, V. Nefedova, D. Quesnel, B. Drach, D. Williams, A. Sim, and A. Shoshani. A date management infrastructure for climate modeling research. In ACM [ACM00], page 116.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Casanova:2000:APS

[COBW00] Henri Casanova, Graziano Obertelli, Francine Berman, and Rich Wolski. The AppLeS Parameter Sweep Template: User-level middleware for the grid. In ACM [ACM00], pages 75–76.

URL http://www.sc2000.org/
proceedings/techpapr/papers/
pap169.pdf.

Chanchio:2000:EPM

[CS00] Kasidit Chanchio and Xian-He Sun. Efficient process migration mechanisms for a heterogeneous distributed computing environment. In ACM [ACM00], page 147. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Chafe:2000:QEA

[CST^{+}00] C. Chafe, S. Shalunov, B. Teitelbaum, M. Gröger, S. Wilson, D. Chisolm, R. Leistikow, and G. Scavone. QoS enabled audio teleportation. In ACM [ACM00], pages 121–122.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Dongarra:2000:TMH

[DDS00] Jack Dongarra, Iain Duff, and Danny Sorensen. Tutorial M7: High-speed numerical linear algebra: Algorithms and research directions. In ACM [ACM00], pages 21–22. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Denneau:2000:BG

[Den00] Monty Denneau. Blue gene.

In ACM [ACM00], page 35.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Drummond:2000:BAT

[Dru00]  Leroy Drummond.  BOF:

The ACTS toolkit development. In ACM [ACM00], page 107.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

deSupinski:2000:BSI

[dS00] Bronis R. de Supinski. BOF: SCICOMP, the IBM SP Scientific Computing User Group. In ACM [ACM00], page 111.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Duke:2000:VV

[DW00] Dennis Duke and Steve Wallach. Venture village. In ACM [ACM00], page 89. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Edelman:2000:LSD

[Ede00] Rich Edelman. Large-scale distributed computing: The benefits and advantages of compute farms. In ACM [ACM00], page 104. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Eigenmann:2000:TMPa

[EM00a] Rudolf Eigenmann and Tim Mattson. Tutorial M6A: Parallel programming with OpenMP: Part I. In ACM [ACM00], page 21. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Eigenmann:2000:TMPb

[EM00b] Rudolf Eigenmann and Tim Mattson. Tutorial M6B: Parallel programming with OpenMP: Part II. In ACM [ACM00], page 23. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Fahey:2000:SCS

[Fah00] Rebecca A. Fahey. STWAVE:

A case study in dual level parallelism. In ACM [ACM00], page 151. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Fortes:2000:PCP

[FEL^{+}00] Jose Fortes, Rudolf Eigenmann, Mark Lundstrom, Valerie Taylor, Miron Livny, Sumalatha Adabala, Renato Figueiredo, Nirav Kapadia, and José Miguel-Alonso. The PUNCH computing portal: Integrating Grid services and Web technologies for high performance computing anytime, anywhere. In ACM [ACM00], page 90. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Ferguson:2000:MLW

[Fer00] Earl Ferguson. MAN/LAN/ WAN — blurring boundaries for switching and routing. In ACM [ACM00], pages 103–104.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Fair:2000:SAA

[FG00] David Fair and Ed Gronke. Software applications accelerate performance with the VI architecture; hot clusters, & what’s next? In ACM [ACM00], pages 36–37.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Fischer:2000:BHS

[Fis00] Markus Fischer. BOF: High speed interconnects: Status and peek into the future. In ACM [ACM00], page 105.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Freisinger:2000:CNG

[Fre00] John Freisinger. The challenge of next generation high performance networks. In ACM [ACM00], page 98. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Feng:2000:FTH

[FT00] W. Feng and P. Tinnakornsrisuphap. The failure of TCP in high-performance computational grids. In ACM [ACM00], page 63. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap174.pdf.

Gamradt:2000:RMS

[Gam00] Derek Gamradt. Remote management of a storage network. In ACM [ACM00], page 100.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Gannon:2000:IE

[GBG^{+}00] Dennis Gannon, Randall Bramley, Madhu Govindaraju, Benjamin Temko, Ken Chiu, and Juan Villacis. The IU Enotebook. In ACM [ACM00], pages 91–92. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Germain:2000:POC

[Ger00]  Robert S. Germain.  Peta-

op computing for large scale biomolecular simulation. In ACM [ACM00], page 34. URL

http://www.sc2000.org/proceedings/
info/fp.pdf.

Grossman:2000:TSI

[GK00] Robert L. Grossman and Vipin Kumar. Tutorial S2: An introduction to high performance data mining. In ACM [ACM00], page 12. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Gropp:2000:PMT

[GKKS00] William D. Gropp, Dinesh K.

Kaushik, David E. Keyes, and Barry F. Smith. Performance modeling and tuning of an unstructured mesh CFD application. In ACM [ACM00], pages 61–62. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap279.pdf.

Gibbons:2000:GSH

[GLP00] T. Gibbons, E. P. Love, and C. Perkins. Gigabit/sec High Definition TV over IP. In ACM [ACM00], page 121.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Gropp:2000:TSU

[GLT00] William Gropp, Ewing (Rusty)

Lusk, and Rajeev S. Thakur. Tutorial S1: Using MPI-2: A tutorial on advanced features of the message-passing interface. In ACM [ACM00], page 11.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Gupta:2000:TSH

[GMM00] Manish Gupta, Samuel P. Midkiff, and Jose E. Moreira. Tutorial S4: High performance numerical computing in Java: Compiler, language, and application solutions. In ACM [ACM00], page 13. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Grossman:2000:PD

[GRC^{+}00] R. Grossman, G. Reinhart, E. Creel, M. Mazzucco, S. Connelly, A. Turinsky, H. Sivakumar, J. Jamiso B. Hollebeek, P. Proropapas, C. Rocke, T. Arons, Y. Guo, S. Hedvall, and P. Milne. Project DataSpace. In ACM [ACM00], pages 116–117. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Green:2000:TWE

[Gre00]  Ron Green.  Tarantella —

Web enable any application on any platform and access from anywhere. In ACM [ACM00], page 95. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Grimshaw:2000:LAP

[Gri00]  Andrew Grimshaw.  Legion

— an applications perspective. In ACM [ACM00], page 99.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Guruswamy:2000:APP

[GRP00] Guru Guruswamy, David Rodriguez, and Mark Postdam. Applications of parallel process hiMAP for large-scale multidisciplinary problems. In ACM [ACM00], page 146.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Govindaraju:2000:RER

[GSC^{+}00] Madhusudhan Govindaraju, Aleksander Slominski, Venkatesh Choppella, Randall Bramley, and Dennis Gannon. Requirements for and evaluation of RMI protocols for scientific computing. In ACM [ACM00], page 76. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap261.pdf.

Gustafson:2000:ANW

[GTH00] John Gustafson, Rajat Todi, and Don Heller. APPMAP: A new way to predict application performance in minutes. In ACM [ACM00], page 146.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Guzenda:2000:ODH

[Guz00] Leon Guzenda. Objectivity/ DB — the high performance database engine. In ACM [ACM00], page 97. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Gilman:2000:MEW

[GZ00] Alfred S. Gilman and Gottfried E. Zimmermann. Middleware and the eSCaped Web. In ACM [ACM00], page 91.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Haerer:2000:AS

[Hae00] Sally Haerer. Awards session. In ACM [ACM00], page 73.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Huffaker:2000:CAS

[HB00a] Bradley Huffaker and Theresa Ott Boisseau. Core AS (Autonomous System) Internet topology. In ACM [ACM00], page 147.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Huskamp:2000:SEP

[HB00b] Jeffrey C. Huskamp and Lisa Bievenue. SC2000 education program. In ACM [ACM00], page 83. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Humphreys:2000:DRS

[HBEH00] Greg Humphreys, Ian Buck, Matthew Eldridge, and Pat Hanrahan. Distributed rendering for scalable displays. In ACM [ACM00], page 60. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap141.pdf.

Hewitt:2000:TMF

[HC00] W. T. Hewitt and I. Curington. Tutorial M3: Framework technologies & methods for large data visualization. In ACM [ACM00], pages 19–20.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Hu:2000:IFG

[HCZ00] Y. Charlie Hu, Alan Cox, and Willy Zwaenepoel. Improving fine-grained irregular shared-memory benchmarks by data reordering. In ACM [ACM00], page 61. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap278.pdf.

He:2000:PAA

[HD00] Yun (Helen) He and Chris H. Q. Ding. Platforms: An accurate arithmetics approach. In ACM [ACM00], page 150.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Hadida:2000:TI

[HDK^{+}00] M. Hadida, A. Durand, Y. Kadobayashi, B. Fink, and M. Ellisman. Telemiscroscopy over IPV6. In ACM [ACM00], page 118.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Henesey:2000:SCI

[Hen00a] Michael J. Henesey. SRC Computers Inc: Background, status, and plans. In ACM [ACM00], page 104. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Henty:2000:PHM

[Hen00b] D. S. Henty. Performance of hybrid message-passing and sharedmemory parallelism for discrete element modeling. In ACM [ACM00], page 50. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap154.pdf.

Hsieh:2000:APE

[HLMR00] Jenwei Hsieh, Tau Leng, Victor Mashayekhi, and Reza Rooholamini. Architectural and performance evaluation of GigaNet and Myrinet interconnects on clusters of small-scale SMP servers. In ACM [ACM00], pages 53–54. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap294.pdf.

Hauser:2000:HCC

[HML^{+}00] Thomas Hauser, Timothy I. Mattox, Raymond P. LeBeau, Henry G. Dietz, and P. George Huang. High-cost CFD on a low-cost cluster. In ACM [ACM00], page 72. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap260.pdf.

Hare:2000:VMI

[HRC^{+}00] Jennifer Hare, Betsy Rice, Jerry Clarke, Margaret Hurley, and William Mattson. Visual MD: An innovative approach to molecular simulation. In ACM [ACM00], pages 152–153. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Hansen:2000:UIW

[HSC00] Scott Hansen, Quinn Snell, and Mark Clement. Utilizing idle workstations in a scheduled parallel computing system. In ACM [ACM00], page 152.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Hoisie:2000:TMP

[HW00] Adolfy Hoisie and Harvey J.

Wasserman. Tutorial M1: Performance analysis and prediction for large-scale scientific applications. In ACM [ACM00], pages 18–19. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Irwin:2000:BWA

[Irw00] Basil Irwin. BOF: Web100:

Automatically tuning TCP for high performance networks. In ACM [ACM00], page 111.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Istrail:2000:FAT

[Ist00] Sorin Istrail. From first assembly towards a new cyberpharmaceutical computing paradigm. In ACM [ACM00], pages 31–32.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Jamison:2000:UJR

[Jam00] John Jamison. Use of Juniper routers in research and education networks. In ACM [ACM00], page 102. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Johnston:2000:CDG

[JGT^{+}00] William E. Johnston, Dennis Gannon, Bill Nitzberg Leigh Ann Tanner, Bill Thigpen, and Alex Woo. Computing and data grids for science and engineering. In ACM [ACM00], pages 70–71. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap253.pdf.

Jones:2000:BPB

[Jon00a] James P. Jones. BOF: Portable Batch System (PBS). In ACM [ACM00], page 111.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Jones:2000:BUA

[Jon00b] Terry Jones. BOF: Using & administering GPFS. In ACM [ACM00], page 109.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Jones:2000:E

[JW00] Stephen Jones and John West. eSCape 2000. In ACM [ACM00], page 89. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Kapadia:2000:ETH

[KAF^{+}00] Nirav H. Kapadia, Sumalatha Adabala, Renato J. Figueiredo, Dolors Royo, José Miguel-Alonso, and Mark S. Lundstrom. Enabling technologies for highperformance computing portals: The PUNCH approach. In ACM [ACM00], pages 147–148.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Kaiser:2000:BPT

[Kai00] Timothy H. Kaiser. BOF: Parallel tools consortium advancing tools for parallel computing. In ACM [ACM00], page 108.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Kaufmann:2000:BTS

[Kau00] Richard Kaufmann. Building a 30 teraOPS supercomputing system. In ACM [ACM00], page 96.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Kearsley:2000:BCR

[Kea00] Simon Kearsley. Basic computational research for drug discovery. In ACM [ACM00], pages 34–35.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Keely:2000:INC

[Kee00]  L. Keely.  Intrepid Net-

work Collaborator (INC). In ACM [ACM00], page 117.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Kiris:2000:PUT

[KKC00] Cetin C. Kiris, Dochan Kwak, and William Chan. Parallel unsteady turbo-pump simulations for liquid rocket engines. In ACM [ACM00], pages 62–63. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap245.pdf.

Kim:2000:RDR

[KLSK00] Seung Jo Kim, Chang Sung Lee, Heon Shin, and Jeong Ho Kim. Rapid design realization of 3D woven composites by HPC. In ACM [ACM00], pages 150–151.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Kramer:2000:S

[Kra00] Bill Kramer. SCinet 2000. In ACM [ACM00], pages 113–115.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Kiefer:2000:HBS

[KS00] Dave Kiefer and Burton Smith. High bandwidth supercomputers of today and tomorrow. In ACM [ACM00], page 97.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Luong:2000:AMB

[LBH00] Phu V. Luong, Clay P. Breshears, and Andy Haas. Advantages of multi-block curvilinear grid and dual-level parallelism in ocean circulation modeling. In ACM [ACM00], page 145.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Lindlan:2000:TFS

[LCM^{+}00] Kathleen A. Lindlan, Janice Cuny, Allen D. Malony, Sameer Shende, Bernd Mohr, Reid Rivenburgh, and Craig Rasmussen. A tool framework for static and dynamic analysis of object-oriented software with templates. In ACM [ACM00], page 68. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap167.pdf.

Lynch:2000:TLP

[LCS00] Vickie E. Lynch, Benjamin A. Carreras, and Nathaniel D. Sizemore. Two levels of parallelism for models of tracer particle transport. In ACM [ACM00], pages 151–152. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Liu:2000:ATS

[LF00] Jian Liu and Jim Ferguson.

Automatic TCP socket buffer tuning. In ACM [ACM00], pages 146–147. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Ludaescher:2000:MBI

[LGMZ00] Bertram Ludaescher, Amarnath Gupta, Maryann E. Martone, and Ilya Zaslavsky. Modelbased integration of heterogeneous neuroscience data sources. In ACM [ACM00], page 149.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Lifka:2000:BWH

[Lif00] David Lifka. BOF: Windows 2000 high peformance computing users. In ACM [ACM00].

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Lindahl:2000:LBE

[Lin00] Greg Lindahl. Linux on the bleeding edge, CPlant and FSL. In ACM [ACM00], page 100.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Ludaescher:2000:TMA

[LM00] Bertram Ludaescher and Richard Marciano. Tutorial M5A: Application building with XML: Standards,tools, and demos: Part I. In ACM [ACM00], pages 20–21.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Lombard:2000:MLC

[Lom00] David Lombard. Moving a large commercial application from SMP to DMP. In ACM [ACM00], pages 41–42. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Li:2000:WGW

[LRSW00] M. Li, O. F. Rana, M. S. Shields, and D. W. Walker. A wrapper generator for wrapping high performance legacy codes as Java/ CORBA components. In ACM [ACM00], page 51. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap135.pdf.

Lan:2000:DLB

[LTB00] Zhiling Lan, Valerie Taylor, and Gregory Bryan. Dynamic load balancing techniques for improving adaptive mesh refinement. In ACM [ACM00], page 147.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Madsen:2000:ICD

[Mad00] Jim Madsen. Internet computing: Distributed computing leads third generation Internet applications. In ACM [ACM00], page 103. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Marczyk:2000:SSB

[Mar00] Jacek Marczyk. Stochastic simulation: Breaking the stagnation and fragmentation of contemporary HPC. In ACM [ACM00], pages 40–41. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Mason:2000:BRE

[Mas00] Ange Mason. BOF: Research experiences for undergraduates program. In ACM [ACM00], page 108. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Mateescu:2000:EPB

[Mat00a] Gabriel Mateescu. Extending the Portable Batch System with preemptive job scheduling. In ACM [ACM00], page 148.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Mattson:2000:CCI

[Mat00b] Tim Mattson. BOF: Cluster computing and the IEEE Task Force on Cluster Computing. In ACM [ACM00], page 109.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Mattson:2000:BOF

[Mat00c] Tim Mattson. BOF: OpenMP and its future developments. In ACM [ACM00], page 106.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

May:2000:TMP

[May00] John M. May. Tutorial M2: Parallel I/O for application developers. In ACM [ACM00], page 19.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Ma:2000:HPV

[MC00] Kwan-Liu Ma and David M. Camp. High performance visualization of time-varying volume data over a wide-area network status. In ACM [ACM00], page 59. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap254.pdf.

McGraw:2000:STP

[McG00] James R. McGraw. SC2000 technical papers/awards. In ACM [ACM00], page 45. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Mesirov:2000:UHG

[Mes00]  Jill P. Mesirov.  Unveiling

the human genome. In ACM [ACM00], page 33. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Makino:2000:TSB

[MFK00] Junichiro Makino, Toshiyuki Fukushige, and Masaki Koga. A 1.349 Tflops simulation of black holes in a galactic center on GRAPE-6. In ACM [ACM00], page 66. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap149.pdf.

Miller:2000:TSP

[MGM00] Barton Miller, Michael Gerndt, and Bernd Mohr. Tutorial S5: Performance analysis and tuning of parallel programs: Resources and tools. In ACM [ACM00], pages 13–14. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Miles:2000:NCF

[Mil00] Doug Miles. New compiler features for high performance computing. In ACM [ACM00], page 103. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Miura:2000:FHP

[Miu00] Kenichi Miura. Fujitsu’s high performance and highly scalable supercomputers and servers. In ACM [ACM00], page 97.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Mitchell:2000:TSM

[MKT00] Scott Mitchell, Patrick Knupp, and Timothy Tautges. Tutorial S6B: Mesh generation for high performance computing. Part II: Mesh generation for massively parallel-based analysis. In ACM [ACM00], page 16. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Marciano:2000:TMA

[ML00] Richard Marciano and Bertram Ludaescher. Tutorial M5B: Application building with XML: Standards,tools, and demos: Part II. In ACM [ACM00], pages 22–23. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

McGuigan:2000:VER

[MM00] Michael D. McGuigan and Stephen Murtagh. Visualization of events from the relativistic heavy ion collider.

In ACM [ACM00], page 153.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Muraki:2000:VCL

[MOK^{+}00] Shigeru Muraki, Masato Ogata, Kagenori Kajihara, Kwan-Liu Ma, and Yutaka Ishikawa. VG cluster: Large scale visual computing system for volumetric simulations. In ACM [ACM00], page 152. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Morris:2000:HPA

[Mor00a] Tom Morris. High performance Alpha clusters. In ACM [ACM00], page 100.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Morton:2000:BPH

[Mor00b] Don Morton. BOF: Promoting high performance computing literacy at low cost. In ACM [ACM00], page 110.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Melody:2000:G

[MS00] S. R. Melody and Jennifer M. Schopf. GridSearcher. In ACM [ACM00], page 149.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Mueller:2000:WWM

[MSB^{+}00] M. Mueller, S. Sanielevici, A. Breckenridge, S. Sekiguchi, J. Brooke, F-P. Lin, and T. Hirayama. World wide metacomputing. In ACM [ACM00], pages 119–120. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Mock:2000:EDS

[MT00] Stephen Mock and Mary Thomas.

eSCape 2000 demonstrations: SDSC GridPort Toolkit applications. In ACM [ACM00], page 90.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Nosrat:2000:NCP

[Nos00] Ramin Nosrat. New capabilities and plans for high performance storage system. In ACM [ACM00], page 99. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Nikolopoulos:2000:DDN

[NPP^{+}00] Dimitrios S. Nikolopoulos, Theodore S.

Papatheodorou, Constantine D. Polychronopoulos, Jesus Labarta, and Eduard Ayguade. Is data distribution necessary in openMP? In ACM [ACM00], page 68. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap192.pdf.

Narumi:2000:TMD

[NSK^{+}00] Tetsu Narumi, Ryutaro Susukita, Takahiro Koishi, Kenji Yasuoka, Hideaki Furusawa, Atsushi Kawai, and Toshikazu Ebisuzaki. 1.34 Tflops molecular dynamics simulation for NaCl with a special-purpose computer: MDM. In ACM [ACM00], page 71. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap218.pdf.

No:2000:IPF

[NT00] Jaechun No and Alok Choudhary Rajeev Thakur. Integrating parallel file I/O and database support for high performance scientific data management. In ACM [ACM00], page 74. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap179.pdf.

Olson:2000:SHR

[OD00] B. Olson and T. Disz. Scalable high-resolution wide area collaboration over the Access Grid. In ACM [ACM00], page 121.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Oehmke:2000:SAA

[OHS00] Robert Oehmke, Janis Hardwick, and Quentin F. Stout. Scalable algorithms for adaptive statistical designs. In ACM [ACM00], page 48. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap194.pdf.

Orman:2000:PTA

[OMS^{+}00] Hilarie Orman, Jamshid Mahdavi, Volker Sander, Wu chun Feng, Stuart Bailey, Lawrence Brakmo, Deepak Bansal, and Brian L. Tierney. Panel: Is TCP an adequate protocol for high performance computing needs? In ACM [ACM00], pages 77–78.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Knupp:2000:TSM

[Owe00] Patrick Knupp Steven J. Owen. Tutorial S6A: Mesh generation for high performance computing. Part I: An overview of unstructured and structured grid generation techniques. In ACM [ACM00], page 14. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Parks:2000:SHB

[Par00] David Parks. Supercomputing — high bandwidth and high performance. In ACM [ACM00], pages 95–96. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Passarelli:2000:DNW

[Pas00] Ben Passarelli. Delivering a new world of flexibility for HPC. In ACM [ACM00], page 101.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Pavlakos:2000:TSC

[PFH^{+}00] Constantine J. Pavlakos, Randall Frank, Patrick Hanrahan, Kai Li, Alan Heirich, and Allen McPherson. Tutorial S10: Commoditybased scalable visualization. In ACM [ACM00], page 18.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Perkins:2000:GSH

[PGM^{+}00] C. Perkins, L. Ghari, A. Mankin, T. Gibbons, and G. Goncher. Gigabit/sec High Definition TV over IP. In ACM [ACM00], page

120. URL http://www.east.
isi.edu/projects/NMAA/.

Pormann:2000:CSC

[PHB^{+}00] John B. Pormann, Craig S. Henriquez, John A. Board, Jr., Donald J. Rose, David M. Harrild, and Alexandra P. Henriquez. Computer simulations of cardiac electrophysiology. In ACM [ACM00], pages 56–57. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap143.pdf.

Plimpton:2000:PAR

[PHBM00] Steve Plimpton, Bruce Hendrickson, Shawn Burns, and Will McLendon III. Parallel algorithms for radiation transport on unstructured grids. In ACM [ACM00], page 57. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap249.pdf.

Perkovic:2000:RSA

[PK00] Dejan Perkovic and Peter J. Keleher. Randomization, speculation, and adaptation in batch schedulers. In ACM [ACM00], page 48. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap238.pdf.

Park:2000:TIW

[PKF^{+}00] Insung Park, Nirav H. Kapadia, Renato J. Figueiredo, Rudolf Eigenmann, and José A. B. Fortes. Towards an integrated,web-executable parallel programming tool environment. In ACM [ACM00], page 49. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap241.pdf.

Pennington:2000:TSD

[PKMB00] Robert Pennington, Patricia Kovatch, Barney Maccabe, and David Bader. Tutorial S3: Design and analysis of high performance clusters. In ACM [ACM00], page 12. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Plesset:2000:CPE

[Ple00]  Erik Plesset.  Concept-to-

production: End to end high performance network monitoring and performance analysis solutions. In ACM [ACM00], page 97.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Parzyszek:2000:GPS

[PNK00] Krzysztof Parzyszek, Jarek Nieplocha, and Ricky A. Kendall. General portable SHMEM library for high performance computing. In ACM [ACM00], pages 148–149. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Pratt:2000:GPS

[PNMB00] T. Pratt, J. Naegle, L. Martinez, and M. Barnaby. Gigabyte per second file transfer in a clustered computing environment. In ACM [ACM00], page 117.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Papka:2000:HRV

[PS00] M. Papka and R. Stevens. Highresolution visualization playback on tiled displays. In ACM [ACM00], page 122.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Pundit:2000:PPA

[PSC^{+}00] Neil Pundit, Marc Snir, Bill Camp, Thomas Sterling, Paul Messina, Rick Stevens, and Pete Beckman. Panel: Petaflops around the corner: When? how is it meaningful? In ACM [ACM00], page 78. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Qiang:2000:SCL

[QRH00] Ji Qiang, Robert D. Ryne, and Salman Habib. Selfconsistent Langevin simulation of Coulomb collisions in chargedparticle beams. In ACM [ACM00], page 58. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap224.pdf.

Ratliff:2000:CCT

[Rat00] Richard Ratliff. Computing challenges in the travel & transportation industry. In ACM [ACM00], page 38. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Roy:2000:MGQ

[RFG^{+}00] Alain J. Roy, Ian Foster, William Gropp, Nicholas Karonis, Volker Sander, and Brian Toonen. MPICH-GQ: Qualityof-service for message passing programs. In ACM [ACM00], page 54. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap234.pdf.

Remondi:2000:SCE

[RH00] Stephen A. Remondi and James Hoch. Satisfying CFD engineering constraints (with parallel processing). In ACM [ACM00], pages 39–40. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Rhoades:2000:HCF

[Rho00]  David Rhoades.  How clus-

ters fill a vendor void. In ACM [ACM00], pages 102–103.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Rivera:2000:TOS

[RT00] Gabriel Rivera and Chau-Wen Tseng. Tiling optimizations for 3D scientific computations. In ACM [ACM00], pages 60–61. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap268.pdf.

Sander:2000:BAS

[San00] Jo Sander. BOF: The ASCI simulation and computer science technology prospectus. In ACM [ACM00], page 106.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Sivakumar:2000:PCA

[SBG00] H. Sivakumar, S. Bailey, and R. L. Grossman. PSockets: The case for applicationlevel network striping for data intensive applications using high speed wide area networks. In ACM [ACM00], pages 63–64. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap240.pdf.

Smarr:2000:PM

[SCF^{+}00] Larry Smarr, Andrew Chien, Ian Foster, Thomas Sterling, David Anderson, and Andrew Grimshaw. Panel: Megacomputers. In ACM [ACM00], page 80.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Smith:2000:OOJ

[SF00] Lance Smith and Rod Fatoohi. An object-oriented job execution environment. In ACM [ACM00], page 49. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap287.pdf.

Schopf:2000:PCG

[SFP^{+}00] Jennifer Schopf, Ian Foster, Cherri Pancake, Marc Snir, and Geoffrey Fox. Panel: Computational grids: A solution looking for a problem? In ACM [ACM00], page 79. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Stout:2000:TSIa

[SJ00a] Quentin F. Stout and Christiane Jablonowski. Tutorial S7A: Introduction to effective parallel computing, Part I. In ACM [ACM00], pages 14–15.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Stout:2000:TSIb

[SJ00b] Quentin F. Stout and Christiane Jablonowski. Tutorial S7B: Introduction to effective parallel computing, Part II. In ACM [ACM00], page 17. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Schroeder:2000:SHG

[SK00] Eleanor Anne Schroeder and James Arthur Kohl. SC2000 HPC games. In ACM [ACM00], page 81. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Schloegel:2000:UAL

[SKK00] Kirk Schloegel, George Karypis, and Vipin Kumar. A unified algorithm for load-balancing adaptive scientific simulations. In ACM [ACM00], page 75.

URL http://www.sc2000.org/
proceedings/techpapr/papers/
pap185.pdf.

Song:2000:MST

[SLJ^{+}00] H. J. Song, X. Liu, D. Jakobsen, R. Bhagwan, X. Zhang, K. Taura, and A. Chien. The MicroGrid: a scientific tool for modeling computational grids. In ACM [ACM00], page 71. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap286.pdf.

Subramanyan:2000:SSB

[SMAF00] Rajesh Subramanyan, José Miguel-Alonso, and José A. B. Fortes. A scalable SNMPbased distributed monitoring system for heterogeneous network computing. In ACM [ACM00], page 52. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap280.pdf.

Sultan:2000:SFT

[SNI00] Florin Sultan, Thu Nguyen, and Liviu Iftode. Scalable fault-tolerant distributed shared memory. In ACM [ACM00], pages 54–55. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap263.pdf.

Sundaram:2000:PSR

[SNT00] Babu Sundaram, Christopher Nebergall, and Steven Tuecke. Policy specification and restricted delegation in Globus proxies. In ACM [ACM00], page 150. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Solchenbach:2000:RWP

[Sol00] Karl Solchenbach. Real-world performance analysis. In ACM [ACM00], page 101.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Spafford:2000:SDI

[Spa00] Eugene H. Spafford. A small dose of infosec. In ACM [ACM00], page 27. URL http://

                                        www.sc2000.org/proceedings/
                                        info/fp.pdf.

Saltz:2000:RSH

[SSK^{+}00a] J. Saltz, A. Sussman, T. Kurc, U. Catalyurik, M. Wheeler, S. Bryant, and M. Peszynska. Reservoir simulation and history matching-grid based computing and interactive dataset exploration. In ACM [ACM00], pages 117–118. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Saltz:2000:TST

[SSK00b] Joel Saltz, Alan Sussman, and Tahsin Kurc. Tutorial S8: Tools and system support for managing and manipulating large scientific datasets. In ACM [ACM00], pages 15–16. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Shan:2000:CTP

[SSOB00] Hongzhang Shan, Jaswinder P.

Singh, Leonid Oliker, and Rupak Biswas. A comparison of three programming models for adaptive applications on the Origin2000. In ACM [ACM00], pages 50–51. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap196.pdf.

Steele:2000:LDT

[Ste00a] Tommy Steele. Liberating design teams via the Internet. In ACM [ACM00], page 96.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Sterling:2000:CCS

[Ste00b] Thomas Sterling. COTS cluster systems for high performance computing. In ACM [ACM00], pages 26–27. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Strohmaier:2000:BTS

[Str00] Eric Strohmaier. BOF: TOP500 supercomputers. In ACM [ACM00], page 109. URL http:/

                                        /www.sc2000.org/proceedings/
                                        info/fp.pdf.

Snell:2000:PPI

[SWCM00] Quinn Snell, Michael Whiting, Mark Clement, and David McLaughlin. Parallel phylogenetic inference. In ACM [ACM00], page 62. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap283.pdf.

Theobald:2000:LCE

[TAK^{+}00] Kevin B. Theobald, Gagan Agrawal, Rishi Kumar, Gerd Heber, Guang R. Gao, Paul Stodghill, and Keshav Pingali. Landing CG on EARTH: A case study of fine-grained multithreading on an evolutionary path. In ACM [ACM00], page 47. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap293.pdf.

Tani:2000:SES

[Tan00] Keiji Tani. Status of the Earth Simulator Project in Japan. In ACM [ACM00], pages 35–36.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Tosic:2000:ALI

[TBW00] Milorad Tosic, Helen Berman, and John Westbrook. Applicationlevel implementation of asynchronous methods in a CORBAbased distributed object database.

In ACM [ACM00], pages 145–

146. URL http://www.sc2000.
org/proceedings/info/fp.pdf.

Terpstra:2000:OSC

[Ter00]  John H. Terpstra.  Open

source community. In ACM [ACM00], page 96. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Schopf:2000:EQS

[Tho00] Thomas Milford,Jennifer M.

Schopf. Extending quality of service on the Grid: Gara and PBS. In ACM [ACM00], page 148.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Traff:2000:IMO

[TRH00] Jesper Larsson Traff, Hubert Ritzdorf, and Rolf Hempel. The implementation of MPI-2 onesided communication for the NEC SX-5. In ACM [ACM00], pages 45–46. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap181.pdf.

Takahashi:2000:PHP

[TSH^{+}00] Toshiyuki Takahashi, Shinji Sumimoto, Atsushi Hori, Hiroshi Harada, and Yutaka Ishikawa. PM2: High performance communication middleware for heterogeneous network environments. In ACM [ACM00], pages 52–53. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap205.pdf.

Tuminaro:2000:PSA

[TT00] Ray S. Tuminaro and Charles Tong. Parallel smoothed aggregation multigrid: Aggregation strategies on massively parallel machines. In ACM [ACM00], page 47. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap232.pdf.

Turek:2000:DBU

[Tur00] David Turek. Deep Blue update. In ACM [ACM00], pages 99–100.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Tierney:2000:TMP

[TWGS00] Brian Tierney, Rich Wolski, Dan Gunter, and Martin Swany. Tutorial M10: Performance tuning and analysis for grid applications. In ACM [ACM00], page 24.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.

Vetter:2000:DST

[VdS00] Jeffrey S. Vetter and Bronis R. de Supinski. Dynamic software testing of MPI applications with Umpire. In ACM [ACM00], page 70. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap208.pdf.

Vadhiyar:2000:ATC

[VFD00] Sathish S. Vadhiyar, Graham E. Fagg, and Jack Dongarra. Automatically tuned collective communications. In ACM [ACM00], page 46. URL http://

                                        www.sc2000.org/proceedings/
                                        techpapr/papers/pap270.pdf.

Voekler:2000:SPC

[Voe00] Goeff Voekler. On the scale and performance of cooperative Web proxy caching. In ACM [ACM00], pages 41–42.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Vazhkudai:2000:SBG

[VT00] Sudharshan S. Vazhkudai and Steven Tuecke. A storage broker for the Globus environment — A ClassAd based implementation. In ACM [ACM00], page 151.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Wallach:2000:KAP

[Wal00] Steven J. Wallach. Keynote address: Petaflops in the year 2009. In ACM [ACM00], page 25.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Wallach:2000:PVC

[WBKG00] Steve Wallach, Matt Blanton, Jackie Kimzey, and Scott Grout. Panel: Venture capital: Who wants to be a billionaire? In ACM [ACM00], page 77.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Wu:2000:TGV

[WBW^{+}00] C. Eric Wu, Anthony Bolmarcich, Marc Snir David Wootton, Farid Parpia, Anthony Chan, Ewing Lusk, and William Gropp. From trace generation to visualization: A performance framework for distributed parallel systems. In ACM [ACM00], pages 69–70. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap228.pdf.

West:2000:TTS

[Wes00] Lynn West. “teaming” technology — staggering performance power through simplicity. In ACM [ACM00], pages 98–99.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Warfield:2000:RTB

[WFG^{+}00] Simon K. Warfield, Matthieu Ferrant, Xavier Gallez, Arya Nabavi, Ferenc A. Jolesz, and Ron Kikinis. Real-time biomechanical simulation of volumetric brain deformation for image guided neurosurgery. In ACM [ACM00], page 56. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap230.pdf.

Wong:2000:ESU

[WOK^{+}00] Adrian T. Wong, Leonid Oliker, William T. C. Kramer, Teresa L. Kaltz, and David H. Bailey. ESP: A system utilization benchmark. In ACM [ACM00], page 52. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap211.pdf.

Wolf:2000:OPI

[Wol00]  Joe H. Wolf.  Optimization

and parallelization with the Intel compilers and the KAI tools. In ACM [ACM00], page 104.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Woolery:2000:III

[Woo00] Robert Woolery. Implementing intelligent infrastructures using SAN appliances. In ACM [ACM00], page 95. URL http://

www.sc2000.org/proceedings/
info/fp.pdf.

Wright:2000:NLN

[Wri00] Margaret H. Wright. Numbers, lots of numbers, and insight, too: Scientific computing 2000. In ACM [ACM00], page 28.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Yang:2000:RFR

[YL00] Chu-Sing Yang and Mon-Yen Luo. Realizing fault resilience in Web-server cluster. In ACM [ACM00], page 55. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap269.pdf.

Zhao:2000:EED

[ZK00] Tao Zhao and Vijay Karamcheti. Expressing and enforcing distributed resource sharing agreements. In ACM [ACM00], page 76. URL http://

www.sc2000.org/proceedings/
techpapr/papers/pap265.pdf.

Zorn:2000:TMC

[ZSS^{+}00] Manfred Zorn, Sylvia Spengler, Horst Simon, Craig A. Stewart, and Inna Dubchak. Tutorial M4: Computational biology and high performance computing. In ACM [ACM00], page 20.

URL http://www.sc2000.org/
proceedings/info/fp.pdf.

Zimmermann:2000:STS

[ZVD00] Gottfried E. Zimmermann, Gregg Vanderheiden, and Dan Deignan. Speech-to-text & speech-to-sign eSCapes hearing barriers. In ACM [ACM00], pages 90–91.

                                        URL http://www.sc2000.org/
                                        proceedings/info/fp.pdf.